the temperature of a body is increased from 27c to 127c,by what fraction would the radiation emitted by it will increase
 

Grade:11

1 Answers

Riddhish Bhalodia
8 years ago
By stephan boltzmann law the radiation power
P \propto T^4
where T is in Kelvin
so
\frac{P_1}{P_2} = \frac{T_1^4}{T_2^4} = \frac{300^4}{400^4} = 81/256